# Boosting In-Context Learning in LLMs Through the Lens of Classical Supervised Learning

### v1

@@ 3.2.1 Affine-logit approximation and leave-subset-out strategy
3.2.2 Context invariance and directional trust region
In the following subsection, we fix the context size $i\in I$ and introduce some enhancement on the proposed method. Note that our surrogate data generation process exposes a well-known limitation of ICL, its sensitivity to the composition and ordering of the context. Specifically, a single query pair $(x,y)$ is evaluated using multiple different sub-contexts $C_{i}$, yielding potentially different logits $\mathbf{m}(x;C_{i})$ and label prediction for the same ground truth label $y$. In essence, an effective calibration method should mitigate this sensitivity, leading to more stable predictions. This motivates incorporating a mechanism to encourage context invariance in the calibrated predictions. To achieve this, we propose augmenting the standard MLE objective (Eq. (6)) with a context-invariance regularization term. Specifically, let $C^{(a)}_{i}$ and $C^{(b)}_{i}$ be any two distinct contexts of size $i$ drawn from $C_{k}$ for evaluating the same query $(x,y)$ in the surrogate data. We aim for the calibrated distributions $\bm{f}(\mathbf{m}(x^{(j)};C^{(a)}_{i});\bm{\theta}^{i})$ and $\bm{f}(\mathbf{m}(x^{(j)};C^{(b)}_{i});\bm{\theta}^{i})$, to be similar. To enforce this similarity, we utilize the symmetric cross-entropy between these two calibrated distributions as a regularizer defined as
$L_{\text{sym}}(\bm{\theta}^{i},x,C^{(a)}_{i},C^{(b)}_{i})$ $=\mathrm{H}\big{(}\bm{f}(\mathbf{m}(x^{(j)};C^{(a)}_{i});\bm{\theta}^{i}),\,\bm{f}(\mathbf{m}(x^{(j)};C^{(b)}_{i});\bm{\theta}^{i})\big{)},$
where $\mathrm{H}(P,Q)\triangleq-\sum_{c=0}^{n-1}(P_{c}\log Q_{c}+Q_{c}\log P_{c})$. This loss term measures the divergence between the two distributions induced by different contexts, penalizing differences in both directions. Then the overall context-invariance penalty is defined by averaging $L_{\text{sym}}$ over all possible pairs of contexts associated with each $x$.
$\text{InvPenalty}(\bm{\theta}^{i})=\sum_{x}\sum_{\{C^{(a)}_{i},C^{(b)}_{i}\}}L_{\text{sym}}(\bm{\theta}^{i},x,C^{(a)}_{i},C^{(b)}_{i}).$ (7)
The full expression of InvPenalty is given in Equation (14) of Appendix D.
On top of ensuring context-invariance, a well-established calibration approach should also take into account the different scenarios induced by the base LLM’s reliability and the size of the context. In particular, strong base LLMs warrant minimal adjustment, while weak ones require more aggressive correction–yet limited examples can mislead both cases, risking overfitting or under-correction.
To balance it, we regularize the calibration by introducing a directional trust region that restricts parameter updates to remain aligned with the base LLM’s logit. Specifically, we constrain the average cosine similarity between each parameter vector $\bm{\theta}^{i}_{c}=[b^{i}_{c},w^{i}_{c}]^{\top}$ and the identity direction $v=[0,1]^{\top}$, which corresponds to the base LLM via
@@ 3.2.2 Context invariance and directional trust region
where $\|\cdot\|_{2}$ refers to $\ell_{2}$-norm and $\tau\in[0,1]$ modulates the trust: large $\tau$ encourages minor scaling adjustments (exploitation), while smaller values permit broader corrections (exploration). This mirrors trust-region principles in policy optimization (e.g., TRPO [24]), adapting model updates based on the confidence in prior predictions.
@@ 3.3 Full algorithm
The final optimization combines this trust constraint with the likelihood loss and a context-invariance regularizer:
$\min_{\bm{\theta}^{i}}\quad$ $\sum_{(\mathbf{m}^{(l)},y^{(l)})\in\mathcal{T}_{i}}-\log f_{y^{(l)}}(\mathbf{m}^{(l)};\bm{\theta}^{i})+\lambda_{\text{inv}}\,\text{InvPenalty}(\bm{\theta}^{i})$
s.t. $\frac{1}{n-1}\sum_{c=1}^{n-1}\frac{(\bm{\theta}_{c}^{i})^{\top}v}{\|\bm{\theta}_{c}^{i}\|_{2}}\geq\tau,$ (8)
where $\lambda_{\text{inv}}>0$ is a hyperparameter controlling the strength of the context-invariance penalty. To solve this optimization problem, we used SciPy’s trust-constr algorithm, a trust-region method designed for constrained optimization. This optimization can be carried out independently for each $i\in I\triangleq\{1,\cdots,k-1\}$, resulting in a set of calibration models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$, each specialized for a particular context length. Additionally, at inference, any sub-context $C_{i}$ can be used to extract logits for a given size $i$. This paves the way for a two-level ensembling strategy to enhance robustness by aggregating predictions across both multiple context lengths and diverse sub-context samples.
Specifically, we train multiple affine-logit models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$ using training sets with different sizes of the context. Then, at inference time, given a test query $x_{\text{test}}$, we first draw $\{C_{i}^{(j)}\}_{j\in\mathcal{M}_{i}}$ from $C_{k}$ for every $i\in I$, where $I$ and $\mathcal{M}_{i}$ are user-defined index sets with size $|\mathcal{M}_{i}|$ and $|I|$. Then we perform intra-size and inter-size ensembling by averaging the calibrated predictions over $\{C_{i}^{(j)}\}_{j\in\mathcal{M}_{i}}$ and across all context sizes $i\in I$ and output the predictive probability of SC for $x_{\mathrm{test}}$ as
$\hat{\mathbf{p}}_{\mathrm{SC}}(x_{\mathrm{test}})=\frac{1}{|I|}\sum_{i\in I}\frac{1}{|\mathcal{M}_{i}|}\sum_{j\in\mathcal{M}_{i}}\bm{f}\bigl{(}\mathbf{m}(x_{\mathrm{test}};C_{i}^{(j)});\hat{\bm{\theta}}^{i}\bigr{)}.$ (9)
The final predicted label is $\hat{y}_{\mathrm{SC}}\in\arg\max_{y_{c}\in\mathcal{Y}}[\hat{\mathbf{p}}_{\mathrm{SC}}]_{c}.$ Overall, this ensembling procedure approximates marginalization over plausible sub-contexts and lengths, significantly improving calibration stability and accuracy. The full algorithm of SC is summarized in Table 2 of Appendix F.
@@ 3.4 Connections to prior work and theoretical insight
In this section, we show the connection of the proposed SC with the existing LM methods and provide a principle approach to theoretically understand these methods from the perspective of supervised learning. Specifically, LM methods rely on one core assumption.
Assumption 1
@@
Under Assumption 1, by a similar derivation in Section 3.2.1, we can show that LM methods are equivalent to assuming
$L^{\ast}_{c}(x)=m_{c}(x;C_{k})+B_{c}(C_{k}),\qquad c=1,\dots,n-1,$ (10)
where $B_{c}(C_{k})=-\log[P_{\text{LLM}}(c|C_{k})/P_{\text{LLM}}(0|C_{k})]$. Therefore, they focus on optimally shifting the decision threshold of the base LLM via estimating $P_{\text{LLM}}(y|C_{k})$, which thus gives an estimator for $B_{c}(C_{k})$. We summarize the existing approaches of estimating $P_{\text{LLM}}(y|C_{k})$ in Table 2 of Appendix D. However, Assumption 1 can be easily violated in practice, causing model mis-specification error.
Instead of imposing Assumption 1, we propose to understand existing LM methods from the perspective of function approximation in the supervised learning. In this case, LM methods basically assume a working model (10). In contrast, the proposed SC considers a strictly larger working model:
@@
This offers a principle framework to compare SC with LM methods and indeed shows that SC generalizes existing LM methods.
Furthermore, within this framework, we analyze these methods via statistical learning theory. Consider a dataset $\mathcal{T}=\{(x^{(j)},y^{(j)})\}_{j=1}^{N}$ of size $N$, and denote by $\hat{f}:=f_{\hat{\bm{\theta}}^{k}}$ the solution minimizing $\mathbb{L}_{k}(\bm{\theta}^{k})$ under $\mathcal{T}$. Let $\mathcal{R}^{*}$ denote the Bayes risk and $\mathcal{R}(\hat{f})$ the $0\text{-}1$ risk of $\hat{f}$. Then, under standard regularity conditions, the excess risk of SC satisfies, with high probability:
$\underbrace{\mathcal{R}(\hat{f})-\mathcal{R}^{\ast}}_{\text{excess risk}}\lesssim\underbrace{\sqrt{D_{\text{KL}}(P^{*}\,\|\,f^{*})-D_{\text{KL}}(P^{*}\,\|\,P^{*})}}_{\text{approximation error}}\;+\;\sqrt{\frac{2(n-1)}{N}},$ (11)
The decomposition leads to the following theoretical insight. Firstly, thanks to the strictly larger working model, SC attains an approximation error that is guaranteed to be no worse than that of LM methods. Secondly, SC estimates $2(n{-}1)$ parameters—one slope and one intercept per non-reference class—while LM methods estimate only $n{-}1$ parameters. This leads to a factor of 2 increase in estimation error, which scales with the number of parameters $d$ as $\mathcal{O}(d)$. This gives LM methods an advantage. However, SC incorporates several variance mitigation strategies to actively control estimation error: (i) diverse training instances generated via Algorithm 1; (ii) explicit regularization through the directional trust region constraint and context invariance penalty; and (iii) ensembling procedure in Algorithm 2. These components work in concert to reduce variance and stabilize learning, enabling the model to fully leverage its lower approximation error.
@@ 4 Experiments and main results
4.1 Experimental settings
Datasets.
We conduct experiments on nine widely used text classification benchmarks spanning sentiment analysis, topic classification, and social media content understanding: SST-2, SST-5 [28], AG News [41], SUBJ [34], TREC [10], Rotten Tomatoes [20], TweetEval-Emotion [18], TweetEval-Hate [1], and Financial PhraseBank [16].
Baselines.
We compare SC against 4 label-marginal-based calibration baselines from prior work, namely Base LLM, CC, BC, and DC. Full implementation details and references are provided in Appendix A.
Models.
We evaluate SC, CC, DC, and BC on three LLMs: Qwen2-7B-Instruct (Qwen) [38], LLaMA-2-7B-Chat-HF (Llama) [30], and Mistral-7B-Instruct-v0.3 (Mistral) [9]. All models are accessed via the Hugging Face API, and no further fine-tuning is applied.
Other Details.
Following prior work, we report Macro-F1 as the primary evaluation metric. Prompts are generated using a fixed template described in Appendix C. We evaluate performance under 4-shot, 8-shot, and 16-shot settings. For each LLM, we run all experiments with 5 random seeds to account for variance due to context sampling. Unless otherwise specified, evaluation is conducted on a held-out test set of 256 examples per dataset.
@@ 4.2 Main results
We report the Macro-F1 performance of five calibration methods across 9 datasets, 5 random seeds, and 3 LLMs (Qwen2-7B-Instruct, LLaMA-2-7B-chat-hf, and Mistral-7B-Instruct-v0.3) under 4-, 8-, and 16-shot settings in Figure 3. Specifically, SC consistently achieves the highest Macro-F1 score across all models and shot counts. Compared to the Base LLM, SC yields improvements of up to +22.6% absolute in Macro-F1 (8-shot on Qwen2-7B-Instruct), and on average provides +11.1% absolute gain across all models and shot configurations. Relative to the strongest competing calibration method (Batch Calibration), SC further improves performance by up to +13.4% (16-shot on Mistral-7B-Instruct-v0.3) and achieves an average gain of +7.1%. Overall, these results confirm that SC offers a robust and generalizable enhancement of Label-Marginal based calibrations in few-shot learning. In addition, our numerical results are aligned with our theory in presented in Section 3.4. As shown in Figure 3, SC achieves the highest average score among all methods due to better approximation error, but also exhibits increased variance in its performance. More detailed numerical results and comparison are given in Appendix G.
Figure 3: Averaged Macro-F1 scores of five methods across 9 datasets, 3 language models, and 5 random seeds under 4-, 8-, and 16-shot settings. Bars represent the mean performance with standard deviations across datasets.
Furthermore, SC delivers a striking improvement on SST-5: in the 8-shot setting with Qwen, it boosts accuracy from 24% (base LLM) and  25% (other methods) to 44%, nearly doubling performance as shown in Figure 4. This substantial gain stems from SC’s unique ability to not just shift logits, but to reverse the decision boundary when necessary as illustrated in Figure 1. For instance, it learns a bias of $-1.29$ and a weight of $-0.19$ for the negative class relative to very negative. This indicates that SC effectively shifts and reorients the LLM’s decision boundary between closely related classes, enhancing overall performance.
@@
We perform ablation studies to understand the contributions of key components within our SC framework. First, we examine the role of the per-class scaling factor. We compare the full SC model with a restricted variant, $\text{SC}^{\ast}$, in which this factor is fixed to 1—so only the bias term is learned—and include additional baseline calibration methods for reference. In our experiments, $\text{SC}^{\ast}$ outperforms the baselines, indicating that estimating an optimal bias under SC framework is more effective than the methods employed by LM approaches. The full SC model achieves still higher performance than $\text{SC}^{\ast}$, suggesting that the flexibility to learn the scaling factor—and therefore to both shift and rescale the LLM’s logits—offers a further advantage. Detailed results are provided in Figures 5 and 6 of Appendix H. Second, we assess whether aggregating calibrators trained with different context sizes improves predictive performance. Specifically, we learn a collection of affine-logit models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$, each fitted on training data whose demonstration context contains $i$ examples. Empirically, both accuracy and Macro-F1 increase monotonically as the size of context $|I|$ grows), indicating that calibrators exposed to heterogeneous amounts of contextual information provide complementary signals that improve robustness and overall predictive quality. Detailed results are provided in Figures  7, 8 and 9 of Appendix H.

### v3

@@ 3.2.1 Affine-logit Approximation and Leave-subset-out Strategy
3.2.2 Context Invariance and Directional Trust-Region
In the following subsection, we fix the context size $i\in I$ and introduce some enhancements to the proposed method. Note that our surrogate data generation process exposes a well-known limitation of ICL, its sensitivity to the composition and ordering of the context. Specifically, a single query pair $(x,y)$ is evaluated using multiple different sub-contexts $C_{i}$, yielding potentially different logits $\mathbf{m}(x;C_{i})$ and label prediction for the same ground truth label $y$. In essence, an effective calibration method should mitigate this sensitivity, leading to more stable predictions. This motivates incorporating a mechanism to encourage context invariance in the calibrated predictions. To achieve this, we propose augmenting the standard MLE objective (Equation (3.2.1)) with a context-invariance regularization term. Specifically, let $C^{(a)}_{i}$ and $C^{(b)}_{i}$ be any two distinct contexts of size $i$ drawn from $C_{k}$ for evaluating the same query $(x,y)$ in the surrogate data. We aim for the calibrated distributions $\bm{f}(\mathbf{m}(x^{(j)};C^{(a)}_{i});\bm{\theta}^{i})$ and $\bm{f}(\mathbf{m}(x^{(j)};C^{(b)}_{i});\bm{\theta}^{i})$, to be similar. To enforce this similarity, we utilize the symmetric cross-entropy between these two calibrated distributions as a regularizer defined as
$L_{\text{sym}}(\bm{\theta}^{i},x,C^{(a)}_{i},C^{(b)}_{i})$
$=\mathrm{H}\big(\bm{f}(\mathbf{m}(x^{(j)};C^{(a)}_{i});\bm{\theta}^{i}),\,\bm{f}(\mathbf{m}(x^{(j)};C^{(b)}_{i});\bm{\theta}^{i})\big),$
where $\mathrm{H}(P,Q)\triangleq-\sum_{c=0}^{n-1}(P_{c}\log Q_{c}+Q_{c}\log P_{c})$. This loss term measures the divergence between the two distributions induced by different contexts, penalizing differences in both directions. Then the overall penalty is defined by averaging $L_{\text{sym}}$ over all possible pairs of contexts associated with each $x$.
$\text{InvPenalty}(\bm{\theta}^{i})=\sum_{\begin{subarray}{c}\{C^{(a)}_{i},C^{(b)}_{i}\}\\ x\end{subarray}}L_{\text{sym}}(\bm{\theta}^{i},x,C^{(a)}_{i},C^{(b)}_{i})$. (6)
The full expression of InvPenalty is given in Equation (12) of Appendix D. On top of ensuring context-invariance, a well-established calibration approach should also take into account the different scenarios induced by the base LLM’s reliability and the size of the context. In particular, strong base LLMs warrant minimal adjustment, while weak ones require more aggressive correction, yet limited examples can mislead both cases, risking overfitting or under-correction. To balance this, we regularize the calibration by introducing a directional trust-region that restricts parameter updates to remain aligned with the base LLM’s logit. Specifically, we constrain the average cosine similarity between each parameter vector $\bm{\theta}^{i}_{c}=[b^{i}_{c},w^{i}_{c}]^{\top}$ and the identity direction $v=[0,1]^{\top}$, which corresponds to the base LLM via
@@ 3.2.2 Context Invariance and Directional Trust-Region
where $\|\cdot\|_{2}$ refers to $\ell_{2}$-norm and $\tau\in[0,1]$ modulates the trust: large $\tau$ encourages minor scaling adjustments (exploitation), while smaller values permit broader corrections (exploration). This mirrors trust-region principles in policy optimization (e.g., TRPO (Schulman et al., 2015)), adapting model updates based on the confidence in prior predictions.
@@ 3.3 Full Algorithm
The final optimization combines this constraint with the likelihood loss and a context-invariance regularizer:
$\min_{\bm{\theta}^{i}}$ $\left\{\sum_{(\mathbf{m}^{(l)},y^{(l)})\in\mathcal{T}_{i}}\hskip-15.00002pt-\log f_{y^{(l)}}(\mathbf{m}^{(l)};\bm{\theta}^{i})+\lambda_{\text{inv}}\text{InvPenalty}(\bm{\theta}^{i})\right\}\;$
$\text{s.t.}\;\frac{1}{n-1}\sum_{c=1}^{n-1}\frac{(\bm{\theta}_{c}^{i})^{\top}v}{\|\bm{\theta}_{c}^{i}\|_{2}}\geq\tau.$ (7)
where $\lambda_{\text{inv}}>0$ is a hyperparameter controlling the strength of the context-invariance penalty. To solve this optimization problem, we used SciPy’s trust-constr algorithm, a trust-region method designed for constrained optimization. This optimization can be carried out independently for each $i\in I\triangleq\{1,\cdots,k-1\}$, resulting in a set of calibration models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$, each specialized for a particular context length. Additionally, at inference, any sub-context $C_{i}$ can be used to extract logits for a given size $i$. This paves the way for a two-level ensembling strategy to enhance robustness by aggregating predictions across both multiple context lengths and diverse sub-context samples. Specifically, we train multiple affine-logit models $\{\hat{\bm{\theta}}^{i}\}_{i\in I}$ using training sets with different sizes of the context. Then, at inference time, given a test query $x_{\text{test}}$, we first draw $\{C_{i}^{(j)}\}_{j\in\mathcal{M}_{i}}$ from $C_{k}$ for every $i\in I$, where $I$ and $\mathcal{M}_{i}$ are user-defined index sets with size $|\mathcal{M}_{i}|$ and $|I|$. Then we perform intra-size and inter-size ensembling by averaging the calibrated predictions over $\{C_{i}^{(j)}\}_{j\in\mathcal{M}_{i}}$ and across all context sizes $i\in I$ and output the predictive probability of SC for $x_{\mathrm{test}}$ as
$\hat{\mathbf{p}}_{\mathrm{SC}}(x_{\mathrm{test}})=\frac{1}{|I|}\sum_{\begin{subarray}{c}i\in I\end{subarray}}\frac{1}{|\mathcal{M}_{i}|}\left(\sum_{j\in\mathcal{M}_{i}}\bm{f}\bigl(\mathbf{m}(x_{\mathrm{test}};C_{i}^{(j)});\hat{\bm{\theta}}^{i}\bigr)\right).$
The final predicted label is $\hat{y}_{\mathrm{SC}}\in\arg\max_{y_{c}\in\mathcal{Y}}[\hat{\mathbf{p}}_{\mathrm{SC}}]_{c}.$ Overall, this ensembling procedure approximates marginalization over plausible sub-contexts and lengths, significantly improving calibration stability and accuracy. The full algorithm of SC is summarized in Algorithm 2 of Appendix E.
@@ 3.4 Connections to Prior Work and Theoretical Insight
In this section, we show the connection of the proposed SC with the existing LM methods and provide a principled approach to theoretically understand these methods from the perspective of supervised learning. Specifically, LM methods rely on one core assumption.
Assumption 1.
@@
Under Assumption 1, the derivation in Section 3.2.1 yields that LM methods are equivalent to assuming
$L^{\ast}_{c}(x)=m_{c}(x;C_{k})+B_{c}(C_{k}),\text{ }c=1,\dots,n-1,$ (8)
where $B_{c}(C_{k})=-\log[P_{\text{LLM}}(c|C_{k})/P_{\text{LLM}}(0|C_{k})]$. Therefore, they focus on optimally shifting the decision threshold of the base LLM via estimating $P_{\text{LLM}}(y|C_{k})$, which thus gives an estimator for $B_{c}(C_{k})$. We summarize the existing approaches of estimating $P_{\text{LLM}}(y|C_{k})$ in Table 2 of Appendix D. However, Assumption 1 can be easily violated in practice, causing model mis-specification error. Therefore, instead of imposing Assumption 1, we propose to understand existing LM methods from the perspective of function approximation in the supervised learning. In this case, LM methods basically assume a working model (8). In contrast, the proposed SC considers a strictly larger working model:
@@
[…]
$\underbrace{\mathcal{R}(\hat{f})-\mathcal{R}^{\ast}}_{\text{excess risk}}$ $\lesssim\underbrace{\sqrt{D_{\text{KL}}(P^{*}\,\|\,f^{*})-D_{\text{KL}}(P^{*}\,\|\,P^{*})}}_{\text{approximation error}}$
$\;+\;\sqrt{\frac{2(n-1)}{N}}.$ (9)
The decomposition leads to the following theoretical insight. Firstly, thanks to the strictly larger working model, SC attains an approximation error that is guaranteed to be no worse than that of LM methods. Secondly, SC estimates 2(n1)2(n{-}1) parameters—one slope and one intercept per non-reference class—while LM methods estimate only n1n{-}1 parameters. This leads to a factor of 2 increase in estimation error, which scales with the number of parameters $d$ as $\mathcal{O}(d)$. This gives LM methods an advantage. However, SC incorporates several variance mitigation strategies to actively control estimation error and fully leverage its lower approximation error: (i) explicit regularization through the directional trust-region constraint and context invariance penalty; and (ii) ensembling procedure in Algorithm 2.
@@ 4 Experiments and Main Results
4.1 Experimental Setup
Datasets. We evaluate our method on nine text classification benchmarks covering sentiment, topic, and social media analysis: SST-2, SST-5 (Socher et al., 2013), AG News (Zhang et al., 2015), SUBJ (Wang and Manning, 2012), TREC (Li and Roth, 2002), Rotten Tomatoes (Pang and Lee, 2005), TweetEval-Emotion (Mohammad et al., 2018), TweetEval-Hate (Basile et al., 2019), and Financial PhraseBank (Malo et al., 2014).
Models and Baselines. We compare SC against the Base LLM and three prior calibration baselines (CC, BC, and DC) on three models: LLaMA-2-7B-Chat-HF (Touvron et al., 2023), Mistral-7B-Instruct-v0.3(Jiang et al., 2023), and Qwen2-7B-Instruct (Yang et al., 2024). All models are used off-the-shelf from Hugging Face without any fine-tuning. Appendix A provides full implementation details for the baselines.
Evaluation. Following prior work, we report Macro-F1 in 4-shot, 8-shot, and 16-shot settings. To ensure robustness, all results are averaged over 5 random seeds on a held-out test set of 256 examples per dataset. Our prompt template is described in Appendix C.
@@ 4.2 Main results
Figure 3 reports the Macro-F1 performance of five calibration methods across our full experimental suite (9 datasets, 3 LLMs, 5 seeds, and 3 few-shot settings). Notably, SC consistently achieves the highest score across all models and shot counts. In particular compared to the Base LLM, SC yields improvements of up to +22.6% absolute in Macro-F1 (8-shot on Qwen2-7B-Instruct), and on average provides +11.1% absolute gain across all models and shot configurations. Relative to the strongest competing calibration method (BC), SC further improves performance by up to +13.4% (16-shot on Mistral-7B-Instruct-v0.3) and achieves an average gain of +7.1%. Overall, these results confirm that SC offers a robust and generalizable enhancement of LM methods in few-shot learning. In addition, our numerical results are aligned with our theory presented in Section 3.4. As shown in Figure 3, SC achieves the highest average score among all methods due to better approximation error, but also exhibits increased variance in its performance. More detailed numerical results and comparison are given in Appendix F.
Figure 3: Average Macro-F1 scores for five methods across 9 datasets and 3 LLMs in 4-, 8-, 16-shots settings. Bars show the mean performance and standard deviation across datasets over 5 random seeds.
Furthermore, SC delivers a striking improvement on SST-5: in the 8-shot setting with Qwen, it boosts accuracy from 24% (base LLM) and  25% (other methods) to 44%, nearly doubling performance as shown in Figure 4. This substantial gain stems from SC’s unique ability to not just shift logits, but to reverse the decision boundary when necessary as illustrated in Figure 1. For instance, it learns a bias of 1.29-1.29 and a weight of 0.19-0.19 for the negative class relative to very negative. This indicates that SC effectively shifts and reorients the LLM’s decision boundary between closely related classes, enhancing overall performance.
@@
To isolate the contributions of our framework’s key components, we conduct several ablation studies which are detailed in Appendix G. First, to isolate the impact of learning the per-class scaling factor $w_{c}$, which underpins SC’s ability to reorient decision boundaries, we compare the full SC model against two alternatives in Figure 5 of Appendix G.1: a restricted variant, $\text{SC}^{*}$ (where $w_{c}$ is fixed to 1, thus only learning an optimal bias term), and other baseline calibration methods. Our experiments reveal that $\text{SC}^{*}$ surpasses these other baselines. This suggests that estimating an optimal bias under SC framework is more effective than methods employed by LM methods. More critically, the full SC model achieves higher performance than $\text{SC}^{\ast}$, suggesting that the flexibility to learn the scaling factor—and therefore to both shift and rescale the LLM’s logits—offers a further advantage.
Second, ensembling across context sizes improves performance. In particular, aggregating calibrators trained at heterogeneous context lengths (increasing $|I|$) yields a consistent improvement in both Accuracy and Macro-F1 across multiple LLMs in the 8-shot setting (Figure 6 of Appendix G.2) and similarly in the 16-shot setting (Figure 7 of Appendix G.2), suggesting that context-size-specific calibrators provide complementary signals.
Third, we investigate the impact of the number of sampled sub-contexts ($m_{i}$) used for prediction averaging within each context-size-specific calibrator during the ensembling phase. In Figure 8 of Appendix G.3, our findings reveal that increasing $m_{i}$ (i.e averaging predictions over a greater number of distinct sub-contexts of size $i$) generally enhances Macro-F1 scores. This suggests that more comprehensive sampling of available context variations for each $i$-shot learner improves the accuracy of the ensemble’s output.
Fourth, we characterize the computational footprint of sub-context (SC) ensembling by reporting wall-clock training time and inference time per 256 test examples. Timing results show that the added compute from SC comes almost entirely at inference time, and it grows approximately linearly with how many sub-context samples we average over. This is because each additional sampled sub-context requires an additional forward pass through the LLM. In contrast, bias-only baselines are essentially insensitive to this choice and stay nearly flat. Concretely, in the 4-shot setting, SC takes 22.91s per 256 test examples when averaging over 1 sampled sub-context, and increases to 134.96s when averaging over 6 sampled sub-contexts, while baselines remain around 10.5s across all settings (Table 10 of Appendix G.4). The same pattern holds for 8-shot: SC increases from 42.83s (1 sampled sub-context) to 260.32s (6 sampled sub-contexts), whereas baselines stay near 11.1s throughout (Table 11 of Appendix G.4).
Fifth, we evaluate the performance contributions of our two main components: the directional trust-region constraint and the context invariance penalty. We observe that removing both constraints improves over the uncalibrated baseline, but adding either the directional trust-region constraint or the context-invariance penalty yields further gains, and combining both achieves the strongest performance (Macro-F1 $0.746$ and Acc. $0.788$; Table 12 of Appendix G.5). This suggests that trust-region and invariance are complementary.
Finally, we assess scalability by repeating our evaluation on the larger LLaMA-13B model. Because these runs are substantially more expensive, we focus on three representative datasets—Rotten Tomatoes, SST-2, and AGNews—and compare against the corresponding 7B variant. The results show that SC scales reliably with model size: on LLaMA-13B, SC remains competitive or achieves the best performance on Rotten Tomatoes and SST-2 (Tables 1314 of Appendix G.6). Moreover, the gains persist and can even become more pronounced as the base model grows. For example, on AGNews, SC improves accuracy from $78.12$ with the 7B model to $88.05$ with the 13B model (Table 15 of Appendix G.6), providing direct evidence that SC’s benefits carry over as underlying model capacity increases.